OpenStudy (anonymous):

In a multiple-choice test, each question has four options. Students will get 2 points for each correct answer; lose 1 point for each incorrect answer; and receive no points for unanswered questions. A student does not know the correct answer for one question. Is it to her advantage or disadvantage to guess an answer? Show your calculations for expected value. Show work.

OpenStudy (anonymous):

expected value right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

suppose you guess. then the probability you get it right is \(\frac{1}{4}\), and the probability you get it wrong is \(\frac{3}{4}\)

OpenStudy (anonymous):

before we compute, it is pretty obvious that it is a bad idea to guess since you only get it right one time out of 4, and will lose 1 point 3 times out of 4, but we can compute the expected value it is what you get, times the probability you get it added up so we compute \[2\times \frac{1}{4}-1\times \frac{3}{4}=\frac{2-3}{4}=-\frac{2}{4}=-\frac{1}{2}\]

OpenStudy (anonymous):

ok that was wrong!!

OpenStudy (anonymous):

\[\frac{2}{4}-\frac{3}{4}=-\frac{1}{4}\]

OpenStudy (anonymous):

in any case it is negative, so you do not want to guess

OpenStudy (anonymous):

easier way is to imagine you answer all questions you get one right for two points , three wrong for -3 points, and your total is then -1 point averaged over the 4 answers you get \(-\frac{1}{4}\) per answer

OpenStudy (anonymous):

in other words, expected value is really a kind of average